Justin has the utility function U = xy, with the marginal utilities MUx = y and MUy = x. The price of x is $2, the price of y is py, and his income is 40. When he maximizes utility subject to his budget constraint, he purchases 5 units of y.

(a)  What must be the price of y and the amount of x consumed?

(b)  Prove that this allocation follows the equi-marginal principle.

(c)  What would be the new bundles of x, y if Px was $3 .


1
Expert's answer
2021-08-31T16:20:15-0400

a)

We know that budget line be

"M=P_XX+P_YY\\\therefore M=40=2X+5Y\\\therefore 40=2X+5Y.............................(1)\\"


Also we know that

"\frac{MU_X}{P_X}=\frac{MU_Y}{P_Y}"


"\therefore \frac{Y}{2}=\frac{X}{5}\\5Y=2X\implies Y=\frac{2}{5}X"


From equation (1)

"40=2x+5(\frac{2}{5})X\\=2X+2X=4X\\X=\frac{40}{4}=10\\\implies X=10 \space units"


Hence

"Y=\frac{2}{5}X=\frac{2}{5}(10)=4"


Y= 4 units


b)

For the equi-marginal principle

"MU_X=P_X\\\therefore Y=P_X"

"Y=4 \space and\space P_X=4"

This proves that allocation above follows the equimarginal principle.


c)

Now, "P_X=3 \space given"

"\frac{Y}{3}=\frac{X}{5}\\5Y=3X\\"

"Y=\frac{3}{5}X"


Again from (1)

"40=2X+5(\frac{3}{5})\\\therefore 40=2X+3X=5X\\\therefore X=\frac{40}{5}=8\implies X=8"


Now

"Y=\frac{3}{5}X=\frac{3}{5}(8)\\\therefore Y=\frac{24}{5}"